OpenStudy (anonymous):

Choose the polar form of the complex number -1 + i.

OpenStudy (anonymous):

@amistre64 @Bahurverse_Luver @cvliboi @dan815 @e.mccormick @freckles @GretaKnows @HazelLuv99 @Micah1mcdugle @rasheem0

OpenStudy (hazelluv99):

idk srry

OpenStudy (anonymous):

its ok

OpenStudy (bahurverse_luver):

I'm no help sorry

OpenStudy (owlcoffee):

So, doing some recheck on the theory, you know that any complex number represents a vector in a complex referential system, so any expression in that form can be represented as: \[ \vec{Z} = a+bi\] so therefore, the module or the "magnitude" of the vector is given by the equation: \[\left| \vec{z} \right|= \sqrt{a ^2 + b^2}\] and the angle it has referential to the x- axis or the "real" component is calculated by: \[arctg(\left| \vec{z} \right|) = \theta \] and that can be used to describe the vector at it´s polar form: \[\left| \vec{z} \right|_{\theta}\]

OpenStudy (owlcoffee):

So, taking the complex number: \[ \vec{z} = -1+i\] we find it's module: \[\left| \vec{z} \right|= \sqrt{(-1)^2 + (1)^2}\] \[\left| \vec{z} \right| = \sqrt{2}\] And now, the angle: You'd have to represent it right now, but I'll do it without a representation and leave that to you. well, the formula was: \[arctg(\left| \vec{z} \right|)= \theta\] So: \[\tan \theta = \left| \vec{z} \right|\] that ould mean: \[Tan \theta = \frac{ 1 }{ -1 } = -1\] so therefore: \[\theta = 135\] and transfering that to polar form: \[\sqrt{2}_{135}\]